A student invested $5000 in two different savings accounts. The first account pays an annual interest rate of 3%. The second account pays an annual interest rate of 4%. At the end of the year, she had earned $185 in interest. How much money did she invest in each account?
Answer by mananth(16949)

3.00% per annum ------------- x
4.00% per annum ------------ y
total investment 5000
Interest----- $185.00
...
x+y=5000 ------------------------1
3.00%x+4.00%y=$185.00
Multiply by 100
3x+4y=$18,500.00 --------2

Multiply (1) by -3
we get
-3x-3y= -15000

Add this to (2)

y=$3,500.00 investment at 4%
Balance $1,500.00 investment at 3%